Q: Suppose there are three types of Apples A, B, and C being sold and consumed. The demand and supply equations for each type are:

DA= 20 - 2PA + 4PB + PC , SA= 4PA - 5

DB=10+3PA - 5PB + 2PC , SB=3PB - 7

DC = 70+4PA +2PB - 5PC , SC=5PC - 16

  • Determine equilibrium prices and quantities.
  • Calculate the elasticity of demand for B with respect to prices of variety A, B and C and interpret the economic meaning of the results
1
Expert's answer
2021-02-22T14:03:10-0500

"-6p_A+4p_B+p_c=-25"

"3p_A-8p_B+2p_C=-17"


"4p_A+2p_B-10p_c=-86"


"p_A=15"


"p_B=12"

"p_C=17"


"q_A=55"


"q_B=29"


"q_C=69"


"E_B=-5\times\frac{12}{29}=-2.07"